A large-scale bakery is laying out a new production process for its packaged bread, which it sells to several grocery chains. It takes 12 minutes to bake the bread. How large of an oven is required so that the company is able to produce 4200 units of bread per hour (measured in the number of units that can be baked simultaneously)?

Consider the baggage check-in process of a small airline. Check-in data indicate that from 9 a.m. to 10 a.m., 240 passengers checked in. Moreover, based on counting the number of passengers waiting in line, airport management found that the average number of passengers waiting for check-in was 30. How long did the average passenger have to wait in line?
Answer #1

Answer:-

1)Given that

time taken to break the bread in the oven =12 minutes

We have to produce a 4200 units of bread per hour.

= 4200/60

= 70 units of bread per minute.

Size of a oven=?

Therefore from little's law inventory (I)=R*T

I= 70*12

I = 840 breads

Therefore the required size of a oven is 840

2) Flow rate (R) = 240 passengers/ 2 hours

R= 120 passengers/hour

Inventory (I) = 30 passengers

According to little's law, T= I/R

T= 30/120

T= 0.25 hours

T = 15 minutes.